OpenStudy (anonymous):

URGENT ASAP find the center of this hyperbola (x+2)^2/49-(y+3)^2/25=1

OpenStudy (anonymous):

it is more like "read" the center looks like \(\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\) a hyperbola with center \((h,k)\)

OpenStudy (anonymous):

so in your case it is \((-2,-3)\)

OpenStudy (anonymous):

Theres supposed to be 2 ordered pairs... idk y tho